The proportion of singers to dancers is 11:6. There are 66 total singers & dancers. How many dancers are there?
Answer by stanbon(75887)

The proportion of singers to dancers is 11:6. There are 66 total singers & dancers. How many dancers are there?
----
11:6 is the same as 11x:6x where dancers are 6x.
-------------
Equation:
11x + 6x = 66
17x = 66
x = 66/17
------------------
6x = 6(66/17) = 23.29.. (# of dancers)
---
Comment: That doesn't sound like the proportion.
====================================================
Cheers,
Stan H.